You have borrowed $100,000 to buy a speed boat. You plan to make monthly equal payments over a 10 year period. The bank has offered you an annual percentage rate (ie nominal rate) of 7% compounded monthly. What is your monthly repayment?
Answer #1
Monthly repayment = Loan amount / Present value of annuity of 1
= $       1,00,000 / 86.12788
= $       1,161.06
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.005833)^-120)/0.005833 i 7%/12 = 0.005833
= 86.12788337 n 10*12 = 120